# FollowIR: Evaluating and Teaching Information Retrieval Models to Follow Instructions

### v1

@@ Abstract
Modern Large Language Models (LLMs) are capable of following long and complex instructions that enable a diverse amount of user tasks. However, despite Information Retrieval (IR) models using LLMs as the backbone of their architectures, nearly all of them still only take queries as input, with no instructions. For the handful of recent models that do take instructions, it’s unclear how they use them. We introduce our dataset FollowIR, which contains a rigorous instruction evaluation benchmark as well as a training set for helping IR models learn to better follow real-world instructions. FollowIR builds off the long history of the TREC conferences: as TREC provides human annotators with instructions (also known as narratives) to determine document relevance, so should IR models be able to understand and decide relevance based on these detailed instructions. Our evaluation benchmark starts with three deeply judged TREC collections and alters the annotator instructions, re-annotating relevant documents. Through this process, we can measure how well IR models follow instructions, through a new pairwise evaluation framework. Our results indicate that existing retrieval models fail to correctly use instructions, using them for basic keywords and struggling to understand long-form information. However, we show that it is possible for IR models to learn to follow complex instructions: our new FollowIR-7B model has significant improvements (over 13%) after fine-tuning on our training set.
@@ 1 Introduction
Large language models (LLMs) have shown impressive results in understanding long complex instructions and have enabled a broad array of tasks (Brown et al., 2020; Achiam et al., 2023). The Information Retrieval (IR) community has adopted the use of LLMs as the main backbone of their neural retrieval architectures and has seen an explosion of progress in semantic search (Karpukhin et al., 2020; Khattab & Zaharia, 2020; Reimers & Gurevych, 2019). However, these existing systems rely on under-specified keywords and short sentences, despite the growing capabilities of LLMs. And even though the retrieval community is increasingly adopting larger IR models for their backbones (such as LLaMA 7B in Pradeep et al. (2023); Ma et al. (2023); Weller et al. (2023)), they are only used to estimate document relevance without any instructions. Moving past this ad-hoc search to search with instructions would enable experts to narrow in on complex information needs using only flexible natural language, adapting to new tasks on the fly. For example, imagine a researcher seeking to identify papers that must contain numerous qualities to be relevant (from a given venue, using a particular method, etc.) while also making sure to avoid conditions that would make the document not-relevant (on the wrong domain, from the wrong time period, etc.).
@@
Recent work has started to move towards search with instructions, but is still understudied, with only a handful of papers on the topic (Muennighoff et al., 2024; Su et al., 2022; Asai et al., 2022). Although these works have shown great initiative in adapting retrieval models to take instructions, these instructions are typically short (less than 10 words), repetitive (only one instruction per dataset e.g. Su et al. (2022); Asai et al. (2022); Wang et al. (2023); Xiao et al. (2023)), and lack evaluation datasets that explicitly measure instruction following – instead using standard retrieval benchmarks.
@@
To address this we introduce FollowIR, which contains a benchmark that explicitly measures the instruction following ability of retrieval models, while also providing training data that includes diverse and realistic instructions. Our key intuition is that many ad-hoc retrieval collections developed as part of TREC (Text REtrieval Conference) conferences provide instructions to human annotators who judge document relevance for a given query, as TREC recognizes that information needs often cannot be specified in only a simple query. These instructions are thorough and complex, including minute details about what makes a document relevant vs not-relevant. Thus if humans can use these TREC instructions to annotate document relevance, so should instruction-following retrieval models (example query and instruction pairs are shown in Figure 1 and Table 1).
@@
We use three deeply-judged111i.e. that a large number of documents have been judged relevant or non-relevant, see Section 2.1 for more. TREC collections as the basis of our evaluation set: TREC Robust 2004 (Voorhees, 2005), TREC Common Core 2017 (Allan et al., 2017), and TREC News 2021 (Soboroff et al., 2020). These collections have been thoroughly annotated in order to evaluate recall in retrieval, frequently with hundreds (or even thousands) of documents judged as relevant or not-relevant. We take the instructions given to human annotators (called narratives) and alter them slightly, re-annotating the relevant documents. We then have paired instructions, which can be used to test how models respond to changing instructions – seeing if they change the set of relevant documents to correspond to the relevance definition in the new altered instructions.
Feature Description Example %
N/A: Query-only
Used for keyword and semantic similarity matching What impact has the Chunnel had on the British economy and/or the life style of the British?
Specifics Detailed specifics on what qualifies as relevant, potentially including or or and conditions Documents discussing the following issues are relevant: - projected and actual impact on the life styles of the British - Long term changes to economic policy and relations - major changes to other transportation systems linked with the Continent … 90.4
Irrelevant Info Not keyword relevant, but helpful for background … The Channel Tunnel, sometimes referred to informally as the Chunnel, is a 50.46-kilometre undersea railway tunnel, opened in 1994 … 34.6
Negation Specifics on what type of documents are not relevant and should be excluded … Documents discussing the following issues are not relevant: - expense and construction schedule - routine marketing ploys by other channel crossers (i.e., schedule changes, price drops, etc.) … 52.9
@@
Table 1: How standard retrieval queries differ from instructions (or narratives): instructions contain more specific details about what is relevant, include less directly-relevant background information, and often have directives about what documents are not relevant, using negation. % indicates the percent of the time these features are present in the original TREC instructions used in FollowIR.
@@
As there are no existing methods to compare pairwise queries in IR, we develop a new evaluation framework to do so, measuring rank-wise score changes (which we call $p$-MRR) of documents given a pair of different instructions with the same query. Results on FollowIR indicate that current models generally fail to follow instructions in retrieval unless they are 3B+ parameters or have not been trained for retrieval. Our analysis shows that these failures are due to two phenomena: the length of the instruction (as models are not used to long instructions) as well as using the instruction for keyword search rather than as an instruction that defines relevance.
@@
Thus, we provide the following contributions: (1) a benchmark for evaluating instruction following in retrieval (FollowIR) consisting of human annotations on top of three already highly-judged corpora, (2) analysis of why current models fail to understand instructions, and (3) training data for teaching retrieval models to follow instructions along with a new open-sourced IR model, FollowIR-7B, that can handle long instructions in IR. [^2] Links to the code, data, and models are available at https://github.com/orionw/FollowIR
Figure 1: A visual depiction of the pairwise evaluation framework: models are evaluated on both the original query and instruction, as well as the altered instruction. If the model correctly understands the instructions, it will change which documents are relevant w.r.t. the alteration (right). Note that the real-world instructions (left) given to TREC annotators includes fine-grained details about what relevance is, as well as instructions containing negation (in bold).
@@ 2.1 TREC Conferences
The United States National Institute of Science and Technology (NIST) created the TREC organization in 1993. Each year TREC sponsors many tracks, or shared tasks, on a given dataset. These tracks range from a variety of topics: anywhere from standard ad-hoc retrieval on news (Soboroff et al., 2018; Soboroff, 2021) to more complex domains such as legal retrieval (Baron et al., 2006; Oard et al., 2008) or retrieval-augmented generation/report-generation (Lawrie et al., 2024).
As part of this process, NIST sponsors annotations for these collections. Typically, this is done by pooling a set of results (runs) from a variety of retrieval models and then annotating them in rank order until funding runs out. To help facilitate annotation, track organizers provide a narrative (or instruction) for each query that will be given to the annotators – however, IR models are only ever given the query. As evaluating total recall would require annotating every document in the collection for every query (which is not feasible for collections with millions of documents), recall error is tested using post-hoc sampling and annotation. Although not every query and document pair can be evaluated, recall for queries is very high. We build off the rigorous evaluation done at TREC by using several of their collections to build FollowIR.
2.2 Instructions for LLMs
Instruction-following LLMs have been popularized by models such as GPT-3 (Brown et al., 2020) and have become a large area of interest for the natural language processing community (Touvron et al., 2023a; Jiang et al., 2023; Groeneveld et al., 2024; Black et al., 2022). There has been much work in evaluating if they can generalize to new instructions (Weller et al., 2020; Wang et al., 2022c; Ouyang et al., 2022), if we can train them to follow instructions without human-annotated data (Wang et al., 2022b; Qin et al., 2023), and applying them to various domains (Zhao et al., 2021; Singhal et al., 2023; Shaghaghian et al., 2020). As the IR community uses LLMs in their models, we seek to broaden the scope of IR to include instructions, aligning it with the broader NLP community.
@@ 2.3 Instructions for Retrieval
Using instructions in retrieval is a nascent area of exploration. Su et al. (2022) and Asai et al. (2022) were two of the earliest works that trained a retrieval model to use instructions along with the query. However, these instructions are typically very short, such as “Retrieve a Wikipedia paragraph that answers this question." Recent work incorporates instructions in smaller models (Xiao et al., 2023; Chen et al., 2023, 2024) as well as others which use Mistral (Jiang et al., 2023) as the backbone of a larger retrieval model that can use instructions: GritLM (Muennighoff et al., 2024) trains Mistral to do both generation and embedding, while Wang et al. (2023) uses Mistral for embeddings only.
Despite this flurry of activity, these efforts do not have an explicit instruction-related retrieval benchmark to evaluate on. Instead, they evaluate on standard retrieval benchmark suites such as MTEB (Muennighoff et al., 2022) and BEIR (Thakur et al., 2021) which do not contain instructions. Thus, these newer instruction-IR models hand-write a few instructions, where typically each instruction is applied to an entire dataset, irrespective of the query. This makes these instructions generic: focused only on the task format, format of the “document" (paragraph, sentence, etc.), and the broad domain. Note that because of this, no current instructions contain any extra background information or negation (Weller et al., 2024) which are commonly found in real-world instructions (Table 1).
In work concurrent to ours, Oh et al. (2024) also propose a dataset to evaluate instructions in retrieval models. Their dataset uses the MSMarco collection, and differs in several crucial aspects: it only has one relevant document per query (e.g. sparsely judged), is GPT-4 generated and validated, focuses on the background of the user (“I am a school teaching looking for …"), and evaluates using the lowest score over $N$ instructions for the same query (measuring robustness). In contrast, we use highly-judged corpora to ensure we can measure recall, use real-world human generated instructions, have human-validated relevance judgements, propose a new paired evaluation protocol, and provide a training dataset and model for teaching instruction-following.
@@ 3 Building FollowIR
We build FollowIR by starting with three existing TREC collections: TREC News 2021 that uses the Washington Post v4 corpus (Soboroff et al., 2020), TREC Robust 2004 using the Disks 4 and 5 collections of various news articles (Voorhees, 2005), and TREC Common Core 2017 using the New York Times Annotated corpus (Allan et al., 2017). Each of these was human annotated to include hundreds of annotations per query (see Table 2), with around 50-180 relevant documents per query on average (and many more not-relevant annotations).
@@
Each of these TREC tracks includes instructions for the human annotators that we use as the instructions. Although using these alone can provide some indication of how well models can follow instructions, it doesn’t explicitly test their instruction following ability. To more carefully isolate this ability in our benchmark, we test whether models can respond to small changes in the instruction.
@@
To accomplish this, we ask two expert human annotators to modify the TREC instructions. However, doing this in a naive way would require re-annotating all the document judgements, a non-trivial task requiring immense annotation efforts. [^3] NIST spends roughly 1-2 million per year: https://trec.nist.gov/pubs/2010.economic.impact.pdf We crucially ask the annotators to make the instructions more specific by including additional aspects to the relevance definition or to include relevance instructions with negation. This decreases the set of documents to re-annotate to only those documents deemed relevant by the original TREC annotators (as it is then a subset of the original relevant documents by definition). This makes the annotation tractable, with only dozens or hundreds of documents to re-annotate per query instead of needed to re-annotate the whole collection (e.g. thousands).
@@
We annotate a subset of the original TREC queries due to cost and overlap: we sample 50 queries from TREC Robust 2004 that do not overlap with TREC Common Core (as Common Core used 50 queries from Robust04 on a new collection), and 30 queries from TREC News 2021. Table 2 shows dataset statistics of judged documents and the final benchmark size. Annotators were asked to change the instructions so that the number of relevant documents was cut in half, thus including a sizeable number of changed relevance judgements. We note that although the number of queries seems small by NLP standards, 30-50 queries is both effective (Webber et al., 2008) and standard in the IR community due to the expense of careful annotation over many documents per query.
@@
Due to differences in retriever quality, if we evaluate by searching over the full collection, each model will retrieve a different number of relevant documents. However, because we evaluate instruction following based on changing the document relevance, models that do poorly in the initial retrieval will have fewer documents which change relevance in the instruction-following evaluation. To rectify this, we instead turn to a reranking task where we include all relevant documents, and use a pool of five models [^4] We use BM25, BGE-base, E5-base-v2, TART-Contriever, and INSTRUCTOR-XL to select the top non-relevant documents. To be able to freely distribute the data due to fair-use laws, we chunk the documents into 400-word passages with 200-word overlap and select the highest scoring passages using MaxP (Dai & Callan, 2019). This enables us to distribute our data, which we do by extending the MTEB evaluation framework (Muennighoff et al., 2022).

### v2

@@ Abstract
Modern Language Models (LMs) are capable of following long and complex instructions that enable a large and diverse set of user requests. While Information Retrieval (IR) models use these LMs as the backbone of their architectures, virtually none of them allow users to provide detailed instructions alongside queries, thus limiting their ability to satisfy complex information needs. In this work, we study the use of instructions in IR systems. First, we introduce our dataset FollowIR, which contains a rigorous instruction evaluation benchmark as well as a training set for helping IR models learn to better follow real-world instructions. FollowIR repurposes detailed instructions—also known as narratives—developed for professional assessors to evaluate retrieval systems. In particular, we build our benchmark from three collections curated for shared tasks at the Text REtrieval Conference (TREC). These collections contains hundreds to thousands of labeled documents per query, making them suitable for our exploration. Through this process, we can measure how well IR models follow instructions, through a new pairwise evaluation framework. Our results indicate that existing retrieval models fail to correctly use instructions, using them for basic keywords and struggling to understand long-form information. However, we show that it is possible for IR models to learn to follow complex instructions: our new FollowIR-7B model has significant improvements after fine-tuning on our training set.
@@ 1 Introduction
Modern language models (LMs) are extensively tuned to be able to follow user instructions faithfully (Chung et al., 2022; Ouyang et al., 2022a; Rafailov et al., 2023; Wang et al., 2023b; Ivison et al., 2023) and safely (Bai et al., 2022; Bianchi et al., 2024). Through these capabilities, LMs are able to successfully tackle a broad range of tasks (Chiang et al., 2024; Liang et al., 2023; Yang et al., 2023; Jimenez et al., 2024; Zeng et al., 2023), even when not explicitly fine-tuned for them.
@@
In contrast to the broader LM community, information retrieval (IR) practitioners and researchers have yet to fully exploit instruction-tuned models. Thanks to their ability to effectively estimate semantic similarity between query and documents, LMs have been adopted as the main backbone of neural retrieval architectures (Karpukhin et al., 2020; Khattab & Zaharia, 2020; Reimers & Gurevych, 2019). However, the vast majority of these systems are fine-tuned to operate exclusively as text spans similarity estimators (Khattab & Zaharia, 2020; Izacard et al., 2021; Nogueira & Cho, 2019; Pradeep et al., 2023; Ma et al., 2023). Moving past these ad-hoc search systems to retrieve with instructions would enable support for complex information needs. For example, imagine a researcher seeking to identify papers that must contain numerous qualities to be relevant (from a given venue, using a particular class of methods, etc.) while also making sure to avoid conditions that would make the document not-relevant (using negative sentiment, using datasets from a given domain, etc.).
@@
Recent work has started to move towards search with instructions, but this topic is still understudied with only a handful of papers (Su et al., 2022; Asai et al., 2022; Muennighoff et al., 2024). In particular, we find their use of instructions to be narrow: instructions are typically short (fewer than 10 words) and repetitive (only one instruction per dataset e.g., Su et al. (2022); Asai et al. (2022); Li & Li (2023); Xiao et al. (2023)). Further, these works lack evaluation datasets that explicitly measure instruction following—instead focusing on standard ad-hoc retrieval benchmarks.
@@
To address these gaps we introduce FollowIR, which consists of (1) a benchmark that explicitly measures the instruction following ability of retrieval models, and (2) training data that includes diverse and realistic instructions. Our key intuition is to leverage instructions developed for professional annotators of IR systems in order to study the capabilities of instruction-following IR models. These instructions are used by annotators to judge document relevance for a given query. Fortunately, the IR field is rich with such data, as these instructions—also known as narratives—are created for all queries in any well-constructed IR dataset. In particular, we use narratives developed for shared tasks at the Text REtrieval Conference111trec.nist.gov (TREC). These instructions are thorough and complex, including minute details about what makes a document relevant vs not-relevant. Thus if annotators can use these TREC instructions to annotate document relevance, so should instruction-following retrieval models (example query and instruction pairs are shown in Figures 1 and 2).
@@
We use three deeply-judged222i.e., that a large number of documents have been judged relevant or non-relevant, see Section 2.1 for more. TREC collections as the basis of our evaluation set: TREC Robust 2004 (Voorhees, 2005), TREC Common Core 2017 (Allan et al., 2017), and TREC News 2021 (Soboroff et al., 2020). These collections have been thoroughly annotated in order to evaluate recall in retrieval, with hundreds to thousands of documents judged as relevant or not-relevant. We take the instructions given to the professional annotators and alter them slightly, manually re-annotating the relevant documents. We then have paired instructions, which can be used to test how models respond to changing instructions; that is, we measure if models update the set of relevant documents to match to the new altered instructions.
@@
Figure 1: How do standard retrieval queries differ from instructions (or narratives)? Instructions contain more specific details about what is relevant, include less directly-relevant background information, and often have directives about what documents are not relevant, using negation. %’s are how often a certain type of content appears in the original TREC instructions used in FollowIR.
@@
As there are no existing methods to compare pairwise queries in IR, we develop a new evaluation framework to do so, measuring rank-wise score changes (which we call $p$-MRR) of documents given a pair of different instructions with the same query. Results on FollowIR indicate that current models generally fail to follow instructions in retrieval unless they are 3B+ parameters or have not been trained for retrieval. Our analysis shows that these failures are due to two phenomena: (1) models are not used to long instructions, and (2) models use instruction for keyword search rather than as a definition of relevance.
@@
In summary, we contribute the following: (1) a benchmark for evaluating instruction following in retrieval (FollowIR) consisting of human annotations on top of three already highly-judged corpora, (2) analysis of why current models fail to understand instructions, and (3) training data for teaching retrieval models to follow instructions along with a new open-sourced IR model, FollowIR-7B, that can handle long instructions in IR. [^3] Links to the code, data, and models are available at https://github.com/orionw/FollowIR
Figure 2: A visual depiction of the pairwise evaluation framework: models are evaluated on the query with the original instruction, and then on the query with the altered instruction. If the model correctly understands the instructions, it will change which documents are relevant w.r.t. the alteration (right). Note that the real-world instructions (left) given to TREC annotators includes fine-grained details about what relevance is, as well as instructions containing negation (in bold).
@@ 2.1 TREC Conferences
The United States National Institute of Science and Technology (NIST) created the TREC organization in 1993. Each year TREC sponsors many tracks, or shared tasks, on a given dataset. These tracks range from a variety of topics: anywhere from standard ad-hoc retrieval on news (Soboroff et al., 2018; Soboroff, 2021) to more complex domains such as legal retrieval (Baron et al., 2006; Oard et al., 2008), or retrieval-augmented generation/report-generation (Lawrie et al., 2024).
As part of this process, NIST sponsors annotations for these collections. Typically, this is done by pooling a set of results (runs) from a variety of retrieval models and then annotating them in rank order until funding runs out. To help facilitate annotation, track organizers provide a narrative (or instruction) for each query that will be given to the annotators—however, IR models are only ever given the query. As evaluating total recall would require annotating every document in the collection for every query (which is not feasible for collections with millions of documents), recall error is tested using post-hoc sampling and annotation. Although not every query and document pair can be evaluated, recall for queries is very high. We build off the rigorous evaluation done at TREC by using several of their collections to build FollowIR.
2.2 Instructions for LMs
Instruction-following LMs have been popularized by models such as InstructGPT (Ouyang et al., 2022a), FLAN (Wei et al., 2022), and T0 (Sanh et al., 2022). They have become a large area of interest for the natural language processing community (Touvron et al., 2023a; Jiang et al., 2023; Groeneveld et al., 2024; Black et al., 2022). There has been much work in evaluating if they can generalize to new instructions (Weller et al., 2020; Wang et al., 2022c; Ouyang et al., 2022b), if we can train them to follow instructions without human-annotated data (Wang et al., 2022b; Qin et al., 2023), and applying them to various domains (Zhao et al., 2021; Singhal et al., 2023; Shaghaghian et al., 2020). As the IR community uses LMs in their pipelines, we seek to broaden the scope of IR to include instructions, aligning it with the broader NLP community.
@@ 2.3 Instructions for Retrieval
Using instructions in retrieval is a nascent area of exploration. Su et al. (2022) and Asai et al. (2022) were two of the earliest works that trained a retrieval model to use instructions along with the query. However, these instructions are typically very short, such as “Retrieve a Wikipedia paragraph that answers this question." Recent work incorporates instructions in smaller models (Xiao et al., 2023; Chen et al., 2023, 2024) as well as others which use Llama (Touvron et al., 2023a; Weller et al., 2023) or Mistral (Jiang et al., 2023) as the backbone of a larger retrieval model that can use instructions: GritLM (Muennighoff et al., 2024) trains Mistral to do both generation and embedding, while Wang et al. (2023a) uses Mistral for embeddings only.
Despite this flurry of activity, these efforts do not have an explicit instruction-related retrieval benchmark to evaluate on. Instead, they evaluate on standard retrieval benchmark suites such as MTEB (Muennighoff et al., 2022) and BEIR (Thakur et al., 2021) which do not contain instructions. Thus, these newer instruction retrieval models hand-write a few instructions, where typically each instruction is applied to an entire dataset, irrespective of the query. This makes these instructions generic: focused only on the task format, format of the “document" (paragraph, sentence, etc.), and the broad domain. Note that because of this, no current instructions contain any extra background information or negation (Weller et al., 2024) which are commonly found in real-world instructions (see Figure 1 for an example of these differences).
In work concurrent to ours, Oh et al. (2024) also propose a dataset to evaluate instructions in retrieval models. Their dataset uses the MS MARCO collection (Nguyen et al., 2016), and differs in several crucial aspects: it only has one relevant document per query (e.g., sparsely judged), is GPT-4 generated and validated, focuses on the background of the user (“I am a school teaching looking for …"), and evaluates using the lowest score over $N$ instructions for the same query (measuring robustness). In contrast, we use highly-judged corpora to ensure we can measure recall, use professionally generated instructions, have human-validated relevance judgements, propose a new paired evaluation protocol, and provide a training dataset and model for teaching instruction-following.
@@ 3 Building FollowIR
We derive FollowIR from three TREC collections: TREC News 2021 (derived from the Washington Post v4 corpus; Soboroff et al., 2020), TREC Robust 2004 (from news articles in Disks 4 and 5 collections; Voorhees, 2005), and TREC Common Core 2017 (from the New York Times Annotated corpus; Allan et al., 2017). Each of these was professionally assessed to include hundreds of annotations per query (see Table 1), with 50-180 relevant documents per query on average (and many more not-relevant annotations).
@@
Each of these TREC tracks includes instructions for the professional annotators that we now also give to the models. Although using these alone can provide some indication of how well models can follow instructions, it doesn’t explicitly test their instruction following ability. To more carefully isolate this in our benchmark, we test whether models can respond to small changes in the instruction.
@@
To accomplish this, we ask two expert annotators to modify the TREC instructions. However, doing this in a naive way would require re-annotating all the document judgements, a non-trivial task requiring immense annotation efforts. [^4] NIST’s budget is \$1–2 million USD/year: trec.nist.gov/pubs/2010.economic.impact.pdf Instead, we task the annotators with making instructions more specific by including additional constraints that narrow the relevance definition. These transformations cause some previously relevant documents to become non-relevant without introducing any new relevant documents from the pool. Therefore, only those documents that were deemed relevant by the original TREC assessors need to be re-annotated. This makes the annotation tractable, with only dozens or hundreds of documents to re-annotate per query instead of a collection of thousands.
@@
We annotate a subset of the original TREC queries due to cost and overlap: we sample 50 queries from TREC Robust 2004 that do not overlap with TREC Common Core (as Common Core used 50 queries from Robust04 on a new collection), and 30 queries from TREC News 2021. Table 1 shows dataset statistics of judged documents and the final benchmark size. Annotators were asked to change the instructions so that the number of relevant documents was cut roughly in half, thus including a sizeable number of changed relevance judgements. We note that although the number of queries seems small by NLP standards, 30-50 queries is both effective (Webber et al., 2008) and standard in the IR community due to the expense of careful annotation over many documents per query.
@@
Due to differences in retriever quality, if we evaluate by searching over the full collection, each model will retrieve a different number of relevant documents. However, because we evaluate instruction following based on changing the document relevance, models that do poorly in the initial retrieval will have fewer documents which change relevance in the instruction-following evaluation. To rectify this, we instead turn to a reranking task where we include all relevant documents, and use a pool of five models [^5] We use BM25, BGE-base, E5-base-v2, TART-Contriever, and INSTRUCTOR-xl. to select the top non-relevant documents. To be able to freely distribute the data due to fair-use laws, we chunk the documents into 400-word passages with 200-word overlap and select the highest scoring passages using MaxP (Dai & Callan, 2019). This enables us to distribute our data, which we do by extending the MTEB evaluation framework (Muennighoff et al., 2022).